### v1

# Developing a Reliable, General-Purpose Hallucination Detection and Mitigation Service: Insights and Lessons Learned

Abstract
Hallucination, a phenomenon where large language models (LLMs) produce output that is factually incorrect or unrelated to the input, is a major challenge for LLM applications that require accuracy and dependability. In this paper, we introduce a reliable and high-speed production system aimed at detecting and rectifying the hallucination issue within LLMs. Our system encompasses named entity recognition (NER), natural language inference (NLI), span-based detection (SBD), and an intricate decision tree-based process to reliably detect a wide range of hallucinations in LLM responses. Furthermore, our team has crafted a rewriting mechanism that maintains an optimal mix of precision, response time, and cost-effectiveness. We detail the core elements of our framework and underscore the paramount challenges tied to response time, availability, and performance metrics, which are crucial for real-world deployment of these technologies. Our extensive evaluation, utilizing offline data and live production traffic, confirms the efficacy of our proposed framework and service.
Developing a Reliable, General-Purpose Hallucination Detection and Mitigation Service: Insights and Lessons Learned
Song Wang, Xun Wang, Jie Mei, Yujia Xie, Sean Muarray, Zhang Li, Lingfeng Wu, Si-Qing Chen , Wayne Xiong Microsoft {sonwang, xunwang, jimei, yujiaxie,}@microsoft.com {murraysean, z67li, lingfw, sqchen, weixi}@microsoft.com
1 Introduction
In the rapidly evolving landscape of natural language processing (NLP), large language models (LLMs) have marked a significant leap forward, unlocking new horizons of capabilities and potentials. However, alongside their remarkable advancements, LLMs bring forth substantial challenges, with "hallucination" standing out as a particularly problematic issue. Hallucination in this context refers to instances when an LLM produces output that is either factually incorrect or not anchored in the supplied input, thus compromising the model’s reliability and the credibility of its applications.
The importance of confronting and mitigating hallucinations in LLM deployments cannot be overstated. It is foundational not only to preserving the integrity of outputs across various applications but also to averting the repercussions of misinformation in sensitive domains. Against this backdrop, our work delves into the nuances of hallucinations within LLMs, placing a special emphasis on intrinsic hallucinations, i.e. errors that can be checked against reference inputs due to their presence in more controlled settings.
A major challenge in hallucination detection and mitigation stems from the complexity and diversity of the phenomenon Ji et al. (2023); Huang et al. (2023); Tonmoy et al. (2024). Hallucinations can vary in their forms and causes depending on the context, model, task, and domain. Therefore, we conducted a comprehensive study to understand the problems by reviewing public benchmark datasets, and also analyzing real hallucinated examples reported by customers and observed during human evaluations. With that, we developed a comprehensive hallucination taxonomy and human labeling guidelines. Furthermore, we developed a GPT4-based detection method, which turns to be very competitive when compared with human annotators.
We aim to develop a production system that provides a general-purpose, fast and accurate hallucination detection and mitigation service with the existing state-of-the-art technologies. To this end, we propose our federated hallucination detection and mitigation framework, as shown in Figure 1, which includes multi-source detection, iterative rewriting and multi-source verification. The multi-source detection module is designed to be diverse and inclusive. It ensembles multiple sources, utilizing named entity recognition (NER), natural languge inference (NLI), span-based-detection (SBD) and other techniques, to recall as much hallucination errors as possible. Based on the AI feedback, we will construct a rewriting prompt dynamically and ask GPT4 to fix the errors with minimal changes. This process can be repeated until there is no hallucination errors or it reaches to a pre-defined rewriting count threshold. Then a multi-source verification module which is similar to the detection module but is optimized for precision, will be applied to further block any remaining hallucinated contents.
After overcoming various challenges, the proposed system has achieved a great balance of accuracy, latency and cost in the production and is being integrated into the customers’ LLM-based applications to reduce and mitigate the escalations about the hallucination issues. Furthermore, we have analyzed the service’s performance on offline benchmark dataset and on the real production traffic, which further demonstrate its effectiveness in serving real-world LLM applications.
In this paper, we detail the key components in our service and highlight critical challenges, lessons and potential limitations, which we believe will benefit future research in this field.
2 Related Works
Hallucination Taxonomy
A widely-adopted classification of hallucination is the intrinsic-extrinsic dichotomyDziri et al. (2021); Huang et al. (2021). Intrinsic hallucination occurs when LLM outputs contradict the provided input, such as prompts. Conversely, extrinsic hallucination occurs when LLM outputs cannot be verified by the information in the input. Recently, reasearchers have proposed more fine-grained taxonomies Pagnoni et al. (2021); Mishra et al. (2024).
We largely followed the categories in Pagnoni et al. (2021) with modifications that reflect the nature and causes of hallucinations in LLM outputs and we also developed a guidelines based on this taxonomy for annotators.
Hallucination Detection
Conventional methods of detecting hallucination can be classified into two types: token-based and sentence-based. The former aims to find hallucinated tokens while the latter is to identify the sentences with hallucinations. Various methods have been developed for identifying hallucinations and most of them leverage the pre-trained LLMs fine tuned on task-specific data Liu et al. (2021); Dziri et al. (2021); Cao et al. (2021); Zha et al. (2023). More recently, LLM with prompt-based methods are also widely used (Manakul et al., 2023; Lei et al., 2023). Though the prompt-based methods require little to no tuning and have competitive performance, they tend to have higher cost and high latency.
In our study, we’ve harnessed the strengths of both methods by employing LLM-based detection for labeling data and creating an ensemble of tailored traditional models such as NER, NLI, and span-based sequence labeling. Our designed detection service conducts a detailed examination of the input text’s semantic and syntactic attributes, allowing it to detect various kinds of hallucinations across different granularities and categories.
Hallucination Mitigation
Hallucination mitigation is to correct the identified hallucinations in the generated responses by LLMs. There are many other ways to reduce hallucination during post-generation. For example RARR Thorne and Vlachos (2021) trained a T5 model using retrieved evidence to generate corrected responses. More recently, researchers have been leveraging LLMs to better utilize hallucination feedback and generate corrections Mündler et al. (2024); Dhuliawala et al. (2023); Lei et al. (2023).
In our work, our rewriter is also LLM-based, leveraging LLM’s self refinement through feedback and reasoning. The key difference is that we have to take into consideration the cost and latency, which demands fewer output tokens, while ensuring the mitigation performance.
Table 1: Hallucination taxonomy and examples. We largely followed the categories in Pagnoni et al. (2021) with modifications that reflect the nature and causes of hallucinations in LLM output generation.
Table 2: Hallucination distribution
3 Hallucination Taxonomy and GPT4-based detection
3.1 Hallucination Taxonomy
It’s important to have a comprehensive understanding of hallucinations and we have dedicated significant effort to developing a hallucination taxonomy by manually analyzing various hallucinated documents.
In Table 1, we listed our taxonomy and examples of hallucinations, and in Table 2, we listed the distribution of different kinds of hallucinations based on human annotations over 34 document summaries. As is shown in the table, hallucinations are of different types and root causes. A majority of them are due to semantic frame errors but discourse errors and pragmatic errors also exist. The variety of hallucinations make the detection a huge challenge. Based on this taxonomy, we define our detailed hallucination guidelines, which improves the inter-annotator agreement in our study.
As stated, we mainly focus on intrinsic hallucinations, which are the errors that can be verified from the source document. While most of the existing hallucination detection solutions are entity-based which treat new entities in the generated summary as hallucinations, from table 2, we found only a small portion,  5% of hallucinations are attributed to new entities. This observation motivates us to develop a federated solution that extends NER based methods to recall more hallucination errors.
3.2 GPT4-based Evaluation
We developed a GPT4-based hallucination detection where we first split the LLM responses into individual sentences and instruct GPT4 to check its factuality sentence by sentence against the source document. This method outperforms on existing methods on multiple datasets (e.g. results on SummAC in Table 4). Other works Lei et al. (2023); Wei et al. (2024) have also developed GPT4-based detection by leveraging LLM’s reasoning capabilities. This GPT4-based detection method plays multiple important roles in our project such as labeling data for training NLI and SBD models and being used in online monitoring of hallucination pipelines in production. However, this method is not suitable to be used in our production system due to its high latency and cost.
3.3 GPT4-based Detection vs Human Annotations
Hallucination annotation is both laborious and costly, requiring clear guidance and extensive background knowledge. We conducted comprehensive studies to understand the gap between human annotation and our designed GPT4 prompt. In table 3, we list our evaluation results based on 20 hallucinated summaries, where GPT4 and humans disagree. We found GPT4 tends to produce higher false positive rates while humans tend to produce high false negative rate, and overall, GPT4 is more accurate than human.
False Positive False Negative Total
Annotators 1 10 11
GPT4 4 5 9
Table 3: Comparison of GPT4 and human
4 Federated hallucination detection and mitigation
In this section, we explain how to develop the components in the proposed federated hallucination detection and mitigation framework as shown in Figure 1.
Figure 1: Federated hallucination detection and mitigation system
4.1 Hallucination Detection
NER-based detection
Named Entity Recognition (NER) aims to identify and categorise key information (entities) in text. By applying NER analysis to the input data, we can spot possible entities that are present in the LLM outouts but not supported by the source document, i.e. hallucinations. We are using an well-known NER service [^1] Azure AI Service: https://azure.microsoft.com/en-us/products/ai-services/ai-language which will return both entity types and the confidences. Figure 2 shows the hallucination distribution over entity types in the generated LLM outputs in a human analysis. Similarly, we conducted the NER analysis on benchmark datasets in target domain, to decide the entity types and confidence thresholds in our NER detection implementation.
Figure 2: Hallucination distribution over entity types
NLI-based detection
Given two input text snippets, a premise and a hypothesis, the task of natural language inference (NLI) aims to predict the relationships: entailment, contradiction or neural. In principle, this is exactly what hallucination detection aims at. However, the premises and hypotheses in most existing NLI datasets (Bowman et al., 2015; Williams et al., 2018; Nie et al., 2020; Schuster et al., 2021) are short (one or two sentences). Therefore, we include a new document-sentence dataset (Kamoi et al., 2023) and use GPT4 to label a new dataset consisting of a diverse set of document-summary pairs. Finally, we finetune the pre-trained encoder of DeBERTa He et al. (2021) with the above datasets, and this model detects hallucinations based on semantic relationship between the document and summary.
Span-based detection
NLI is a sentence level hallucination detection while NER is restricted to a predefined set of named entities, and we further explore more fine-grained hallucination detection. Specifically, we train a token level hallucination detection model to provide more fine-grained AI feedback, e.g. hallucinated text spans. We start with the dataset labeled for NLI model and further ask GPT4 to highlight the hallucinated text spans when it is hallucinated. As shown in Figure 3, we start with a pre-trained replaced token detection head from DeBERTa He et al. (2021), and adapt it with GPT-4 generated data like a sequence labeling task. We refer to this model as Span-based detection model (SBD).
Figure 3: Span based hallucination detection
Multi-source ensemble
We believe the combination of multi-sources is necessary to further boost performance with minimal cost. For this part, we collected about 10k training examples from the target traffic, where the label (hallucinated vs not-hallucinated) is produced by GPT4 detection and the features are outputs from NER, NLI and SBD detection methods. We adopt the Gradient boosting decision tree to leverage the diverse AI feedback and finetune a model using scikit-learn to generate a single numerical value indicating whether the response or summary is hallucinated.
4.2 GPT4-based rewriting
With AI feedback from different sources, we compose a prompt based on a pre-defined template and ask GPT4 to fix it. Initially, we devise prompt version v1 that does an exhaustive chain-of-thoughts reasoning and does a full rewriting, which gives better performance but higher latency. Later on, to meet the latency requirement in the product, we developed rewriting prompt v2, which reduces chain-of-thought reasoning and then skip the rewriting if not hallucinated or otherwise only rewrite the hallucinated sentences instead of the full response.
5 Main Results
5.1 Detection Results
We compare the different detection methods’ performance on both internal and public benchmark datset. As shown in Table 4, NER tends to have lower recall but relatively higher precision in public SummAC Laban et al. (2022). It’s notable that of the three tailored models, the SBD method outperforms the rest in all metrics, showcasing the effectiveness of detailed detection - an insight that is valuable to highlight.
Table 4: Performance of different methods on the internal benchmark and public SummAC.
5.2 Mitigation Results
We define mitigation rate as percentage of hallucination being fixed and use GPT4 output tokens length as proxy to the latency and cost. As shown in Table 5, prompt v2 is more token-efficient and achieves good balance of performance, latency and cost.
Table 5: Rewriting to balance the quality and latency.
5.3 Performance in Production
As mentioned above, we have developed two pipelines: 1) the detection-only pipeline and 2) the detection with mitigation pipeline. They are integrated into LLM-based products to mitigate the customer’s complaints about hallucination. In practice, we start with the detection-only pipeline and block the hallucinated context from affecting the customers. Gradually, we moved on to the full pipeline of detection and mitigation in the user experience as shown in Figure 4.
Figure 4: User experience of Hallucination Detection & Mitigation Pipeline in Production
Accurately measuring the effectiveness of hallucination detection and mitigation in a production environment is a complex task. We adopt two approaches: we conduct offline human evaluation on some benchmark datasets and we develop a mirror traffic-based flighting system that leverages GPT4 for online quality measurement. This online system applies mirrored traffic to various pipelines and uses GPT4 to assess hallucination rates and the overall quality of rewritten content, and it better monitors the dynamic nature of real-world applications.
The human evaluation result in Table 6 shows that our detection pipelines are very effective. We experimented on key point generation datasets which contain 469 and 161 documents separately. On both datasets, the proposed methods are able to detect hallucinations with precision of above 50% and reduces more than half of the detected hallucinations with precision of above 60%.
Table 6: Offline human evaluation of detection and mitigation.
When applied to a real product scenario, the precision cannot be evaluated directly as we cannot check customer data. As shown by the online result in Table 7, the GPT4-based mitigation can greatly reduce the hallucinated responses and thus the block rate at expense of acceptable increase of latency.
Table 7: Online monitoring and comparisons of the pipelines based on flighting traffic.
6 Challenges, Learnings & Limitations
We have encountered a number of considerable challenges when we integrate this federated hallucination detection and mitigation service hallucination into real products. We have suggested various workable solutions, although some are imperfect, and we outline them with the aim of assisting the community in acknowledging and addressing these challenges collaboratively.
6.1 Measurement of Effectiveness in Production
Accurately measuring the effectiveness of hallucination detection and mitigation and the values bringing to customers in a production environment is very challenging. We have designed a system which applies mirrored traffic to various pipelines and uses GPT4 to assess hallucination rates and the overall quality of rewritten content. However, the GPT4-based measurement has limitation and ensuring the reliability of these measurements and their alignment with human judgment remains an ongoing challenge, necessitating continuous refinement and validation.
6.2 Handling Multilingual and Long Source Documents
We have add several major non-English training data in NLI and SBD models to support the multilingual use case and using segmentation based approach to deal with long source documents. However, managing multilingual even cross-lingual inputs and long source documents adds another layer of complexity. Also it is still an open research problems how to do effective modeling of the long source document.
6.3 Deep Customization Needs for Various Scenarios
Different usages and user circumstances call for tailored adjustments. In the production, we’ve calibrated our ensemble-based detection for greater accuracy with a reduced block rate, while also adjusting our rewriting for decreased latency at the cost of some mitigation power. As we move forward, incorporating additional domain or task-specific datasets may enhance our detection capabilities. Concurrently, researching novel detection techniques or focusing exclusively on detecting certain types or severe cases of hallucinations might be more practical for real-world applications. Furthermore, we should consider investigating more efficient mitigation strategies involving smaller language models.
7 Conclusions
To conclude, we developed a reliable and high-speed production service to detect and mitigate intrinsic or close-world hallucinations in LLMs. Our service combines named entity recognition (NER), natural language inference (NLI), span-based detection (SBD) models, and a decision tree analysis to identify diverse set of hallucinations with high accuracy. We also develop an efficient rewriting mechanism to ensure the quality and fidelity of LLM outputs.
Hallucination is a key problem facing by LLM applications. Our journey from idea to deployment revealed several challenges in delivering high-quality solutions under real-world constraints. We had to balance latency, availability, and quality in production settings.
Our work has broad applicability and potential for future impact, as LLMs are used in diverse sectors that require robust hallucination verification mechanisms. Our service can adapt to different LLM applications, each with different challenges and demands. A key next step is to address extrinsic or open-world hallucinations, which require integrating external knowledge bases and information verification systems.
8 Ethical Considerations
Ethical considerations are paramount in the development and deployment of hallucination mitigation systems. Ensuring transparency in detection and mitigation processes, providing clear explanations for decisions, and safeguarding user data are essential components of our ethical framework. Balancing these ethical imperatives with technical and operational demands is a complex but necessary challenge.
Appendix A Appendix
This is an appendix.

### v2

@@
-Developing a Reliable, General-Purpose Hallucination Detection and Mitigation Service: Insights and Lessons Learned
+Developing a Reliable, Fast, General-Purpose Hallucination Detection and Mitigation Service
 Abstract
@@
 Our system encompasses named entity recognition (NER), natural language inference (NLI), span-based detection (SBD), and an intricate decision tree-based process to reliably detect a wide range of hallucinations in LLM responses.
-Furthermore, our team has crafted a rewriting mechanism that maintains an optimal mix of precision, response time, and cost-effectiveness.
+Furthermore, we have crafted a rewriting mechanism that maintains an optimal mix of precision, response time, and cost-effectiveness.
 We detail the core elements of our framework and underscore the paramount challenges tied to response time, availability, and performance metrics, which are crucial for real-world deployment of these technologies.
 Our extensive evaluation, utilizing offline data and live production traffic, confirms the efficacy of our proposed framework and service.
-Developing a Reliable, General-Purpose Hallucination Detection and Mitigation Service: Insights and Lessons Learned
+Developing a Reliable, Fast, General-Purpose Hallucination Detection and Mitigation Service
 Song Wang, Xun Wang, Jie Mei, Yujia Xie, Sean Muarray, Zhang Li, Lingfeng Wu, Si-Qing Chen, Wayne Xiong Microsoft {sonwang, xunwang, jimei, yujiaxie,}@microsoft.com {murraysean, z67li, lingfw, sqchen, weixi}@microsoft.com
@@
 Hallucination in this context refers to instances when an LLM produces output that is either factually incorrect or not anchored in the supplied input, thus compromising the model’s reliability and the credibility of its applications.
-The importance of confronting and mitigating hallucinations in LLM deployments cannot be overstated.
-It is foundational not only to preserving the integrity of outputs across various applications but also to averting the repercussions of misinformation in sensitive domains.
-Against this backdrop, our work delves into the nuances of hallucinations within LLMs, placing a special emphasis on intrinsic hallucinations, i.e.
-errors that can be checked against reference inputs due to their presence in more controlled settings.
-A major challenge in hallucination detection and mitigation stems from the complexity and diversity of the phenomenon Ji et al.
-(2023); Huang et al.
-(2023); Tonmoy et al.
-(2024).
-Hallucinations can vary in their forms and causes depending on the context, model, task, and domain.
-Therefore, we conducted a comprehensive study to understand the problems by reviewing public benchmark datasets, and also analyzing real hallucinated examples reported by customers and observed during human evaluations.
-With that, we developed a comprehensive hallucination taxonomy and human labeling guidelines.
-Furthermore, we developed a GPT4-based detection method, which turns to be very competitive when compared with human annotators.
-We aim to develop a production system that provides a general-purpose, fast and accurate hallucination detection and mitigation service with the existing state-of-the-art technologies.
-To this end, we propose our federated hallucination detection and mitigation framework, as shown in Figure 1, which includes multi-source detection, iterative rewriting and multi-source verification.
-The multi-source detection module is designed to be diverse and inclusive.
-It ensembles multiple sources, utilizing named entity recognition (NER), natural languge inference (NLI), span-based-detection (SBD) and other techniques, to recall as much hallucination errors as possible.
-Based on the AI feedback, we will construct a rewriting prompt dynamically and ask GPT4 to fix the errors with minimal changes.
-This process can be repeated until there is no hallucination errors or it reaches to a pre-defined rewriting count threshold.
-Then a multi-source verification module which is similar to the detection module but is optimized for precision, will be applied to further block any remaining hallucinated contents.
-After overcoming various challenges, the proposed system has achieved a great balance of accuracy, latency and cost in the production and is being integrated into the customers’ LLM-based applications to reduce and mitigate the escalations about the hallucination issues.
-Furthermore, we have analyzed the service’s performance on offline benchmark dataset and on the real production traffic, which further demonstrate its effectiveness in serving real-world LLM applications.
-In this paper, we detail the key components in our service and highlight critical challenges, lessons and potential limitations, which we believe will benefit future research in this field.
-2 Related Works
+Therefore, the importance of confronting and mitigating hallucinations in LLM deployments cannot be overstated.
+Detecting and mitigating hallucinations present tough challenges, actively explored in recent research as evidenced by several survey papers (Ji et al., 2023; Huang et al., 2023; Tonmoy et al., 2024).
+There are also different levels of hallucinations spanning from minor inconsistencies to blatant fabrications, and they can have different effects for different applications and users.
+Against this backdrop, our work delves into the nuances of hallucinations within LLMs, placing a special emphasis on intrinsic hallucinations i.e.
+errors that can be checked against reference inputs.
+Developing a general-purpose, fast and accurate hallucination detection and mitigation service is an extremely difficult task given the existing state-of-the-art technologies.
+To this end, we present a pragmatic solution as shown in Figure 1, which includes three modules: multi-source detection, iterative rewriting and multi-source verification.
+We will discuss the components in details in Section 4.
+Figure 1: End-2-end hallucination detection and mitigation system
+Our contributions are as follows.
+First, we present a novel detection system capable of detecting different types of hallucinations with high accuracy.
+The system operates in real-time (low latency) and is suitable for large-scale applications (low cost).
+This approach leverages multiple hallucination detection methods—including named entity recognition (NER), natural language inference (NLI), and span-based detection—and ensembles multiple AI feedbacks using Gradient Boosting Decision Trees (GBDT).
+Second, we propose a rewriting system for hallucination removal utilizing large language models (LLMs).
+After testing various strategies, we developed an effective rewriting solution that balances quality and latency.
+Third, we conducted comprehensive experiments, analyses, and evaluations, demonstrating that our methods are effective and providing insights valuable for other researchers and industry scientists.
+The results are convincing and highlight the applicability of our methods in real-world scenarios.
+2 Related Work
 Hallucination Taxonomy
@@
 More recently, LLM with prompt-based methods are also widely used (Manakul et al., 2023; Lei et al., 2023).
-Though the prompt-based methods require little to no tuning and have competitive performance, they tend to have higher cost and high latency.
+Though the prompt-based methods require little to no tuning and have competitive performance, they tend to have higher cost and higher latency.
 In our study, we’ve harnessed the strengths of both methods by employing LLM-based detection for labeling data and creating an ensemble of tailored traditional models such as NER, NLI, and span-based sequence labeling.
@@
 The key difference is that we have to take into consideration the cost and latency, which demands fewer output tokens, while ensuring the mitigation performance.
-Table 1: Hallucination taxonomy and examples.
-We largely followed the categories in Pagnoni et al.
-(2021) with modifications that reflect the nature and causes of hallucinations in LLM output generation.
+Table 1: Hallucination taxonomy and examples. We largely followed the categories in Pagnoni et al. (2021) with modifications that reflect the nature and causes of hallucinations in LLM outputs.
 Table 2: Hallucination distribution
@@
 3.1 Hallucination Taxonomy
-It’s important to have a comprehensive understanding of hallucinations and we have dedicated significant effort to developing a hallucination taxonomy by manually analyzing various hallucinated documents.
-In Table 1, we listed our taxonomy and examples of hallucinations, and in Table 2, we listed the distribution of different kinds of hallucinations based on human annotations over 34 document summaries.
-As is shown in the table, hallucinations are of different types and root causes.
-A majority of them are due to semantic frame errors but discourse errors and pragmatic errors also exist.
-The variety of hallucinations make the detection a huge challenge.
-Based on this taxonomy, we define our detailed hallucination guidelines, which improves the inter-annotator agreement in our study.
-As stated, we mainly focus on intrinsic hallucinations, which are the errors that can be verified from the source document.
-While most of the existing hallucination detection solutions are entity-based which treat new entities in the generated summary as hallucinations, from table 2, we found only a small portion,  5% of hallucinations are attributed to new entities.
-This observation motivates us to develop a federated solution that extends NER based methods to recall more hallucination errors.
-3.2 GPT4-based Evaluation
-We developed a GPT4-based hallucination detection where we first split the LLM responses into individual sentences and instruct GPT4 to check its factuality sentence by sentence against the source document.
-This method outperforms on existing methods on multiple datasets (e.g.
-results on SummAC in Table 4).
-Other works Lei et al.
+We started with developing a hallucination taxonomy by manually analyzing various hallucinated model outputs, mainly from some internal summarization systems, which is based on a state-of-the-art encoder-decoder model about 1B parameters.
+We randomly collected 500 samples from production systems and benchmark systems, including ChatGPT, and manually identified the 34 hallucinated outputs (two of the authors).
+In Table 1, we list our taxonomy and examples of hallucinations, which show that hallucination is of different types and root causes.
+Most of them are due to semantic frame errors, but discourse errors and pragmatic errors also exist.
+In Table 2, we list the distribution of different types of hallucinations.
+As stated above, we focus mainly on intrinsic hallucinations, which are errors that can be verified from the source document.
+Although most of the existing hallucination detection solutions are entity-based which treat new entities in the generated summary as hallucinations, from table 2, we found only a small portion,  5% of intrinsic hallucinations are attributable to new entities.
+This observation motivates us to develop an ensemble-based solution that extends NER-based methods to recall more hallucination errors.
+3.2 GPT4-based Detection
+We developed a GPT4-based hallucination detection as follows (we cannot share the prompt due to proprietary limitations): First, we transform the LLM outputs so that individual sentences are placed on separate lines.
+We then instruct GPT-4 to evaluate each sentence by comparing it against the source document and provide reasons if a sentence is determined to be a hallucination.
+If any sentence within the output is identified as hallucinated, the entire output is considered hallucinated.
+This approach utilizes the Chain of Thought (CoT) technique, and has been shown to outperform existing methods on multiple datasets; for instance, see the results on SummAC in Table 4.
+Other studies Lei et al.
 (2023); Wei et al.
-(2024) have also developed GPT4-based detection by leveraging LLM’s reasoning capabilities.
-This GPT4-based detection method plays multiple important roles in our project such as labeling data for training NLI and SBD models and being used in online monitoring of hallucination pipelines in production.
-However, this method is not suitable to be used in our production system due to its high latency and cost.
-3.3 GPT4-based Detection vs Human Annotations
-Hallucination annotation is both laborious and costly, requiring clear guidance and extensive background knowledge.
-We conducted comprehensive studies to understand the gap between human annotation and our designed GPT4 prompt.
-In table 3, we list our evaluation results based on 20 hallucinated summaries, where GPT4 and humans disagree.
-We found GPT4 tends to produce higher false positive rates while humans tend to produce high false negative rate, and overall, GPT4 is more accurate than human.
+(2024) have also demonstrated the effectiveness of leveraging GPT-4’s reasoning capabilities for hallucination detection.
+Additionally, we examined discrepancies between human annotation and our GPT-4-based evaluation using 20 model outputs from a benchmark dataset of 1,400 samples (see Table 3).
+Our analysis indicates that while GPT-4 tends to have higher false positive rates, human annotators often show higher false negative rates.
+Nonetheless, GPT-4’s labeling is comparable to human efforts in overall error counts and can enhance annotator productivity by combining the model’s high recall with human precision.
+However, the study’s limited sample size and the specific design of prompts may constrain the generalizability of these findings, indicating a need for further research.
 False Positive False Negative Total
@@
 GPT4 4 5 9
-Table 3: Comparison of GPT4 and human
-4 Federated hallucination detection and mitigation
-In this section, we explain how to develop the components in the proposed federated hallucination detection and mitigation framework as shown in Figure 1.
-Figure 1: Federated hallucination detection and mitigation system
-4.1 Hallucination Detection
+Table 3: Error analysis of the inconsistency of GPT4 and human,
+4 Hallucination Detection and Mitigation
+In this section, we introduce the hallucination checking method as shown in Figure 1 tailored for intrinsic hallucinations, which employes an ensemble method that leverages multiple techniques—including named entity recognition (NER), natural language inference (NLI), and sequence labeling—to detect inaccuracies.
+None of these models are LLM-based mainly for two main reasons: First, we aim to achieve real-time detection with high accuracy suitable for large-scale applications.
+The substantial costs associated with calling LLM models are not feasible.
+Second, LLMs are not effective at detecting hallucinations in their own outputs.
+Previous works have shown that LLMs tend to believe their outputs are correct and are difficult to persuade otherwise Farquhar et al.
+(2024); Quevedo et al.
+(2024).
+4.1 Hallucination detection methods
 NER-based detection
-Named Entity Recognition (NER) aims to identify and categorise key information (entities) in text.
-By applying NER analysis to the input data, we can spot possible entities that are present in the LLM outouts but not supported by the source document, i.e.
-hallucinations.
-We are using an well-known NER service [^1] Azure AI Service: https://azure.microsoft.com/en-us/products/ai-services/ai-language which will return both entity types and the confidences.
-Figure 2 shows the hallucination distribution over entity types in the generated LLM outputs in a human analysis.
-Similarly, we conducted the NER analysis on benchmark datasets in target domain, to decide the entity types and confidence thresholds in our NER detection implementation.
+Named Entity Recognition (NER) aims to identify and categorize key information (entities) in text.
+By applying NER analysis to the input data, we can spot possible entities that are present in the LLM outputs but not supported by the source document—that is, hallucinations.
+We are using a well-known NER service [^1] Azure AI Service: https://azure.microsoft.com/en-us/products/ai-services/ai-language which returns both entity types and their confidence scores.
+We apply this NER service to detect hallucinations, and Figure 2 shows the common entity types among the hallucinated LLM outputs.
+Additionally, we conducted NER analysis on benchmark datasets in the target domain to determine the entity types and confidence thresholds for our NER detection implementation.
 Figure 2: Hallucination distribution over entity types
 NLI-based detection
-Given two input text snippets, a premise and a hypothesis, the task of natural language inference (NLI) aims to predict the relationships: entailment, contradiction or neural.
-In principle, this is exactly what hallucination detection aims at.
-However, the premises and hypotheses in most existing NLI datasets (Bowman et al., 2015; Williams et al., 2018; Nie et al., 2020; Schuster et al., 2021) are short (one or two sentences).
-Therefore, we include a new document-sentence dataset (Kamoi et al., 2023) and use GPT4 to label a new dataset consisting of a diverse set of document-summary pairs.
-Finally, we finetune the pre-trained encoder of DeBERTa He et al.
-(2021) with the above datasets, and this model detects hallucinations based on semantic relationship between the document and summary.
+In natural language inference (NLI), given two input text snippets—a premise and a hypothesis—the task is to predict their relationship: entailment, contradiction, or neutral.
+In principle, this aligns with the goal of hallucination detection.
+However, in most existing NLI datasets (Bowman et al., 2015; Williams et al., 2018; Nie et al., 2020; Schuster et al., 2021), the premises and hypotheses are short (one or two sentences).
+To address this, we included a new document-sentence dataset (Kamoi et al., 2023) and used GPT-4, as described in Section 3.2, to label a diverse set of document-summary pairs from various public and internal sources.
+Finally, we fine-tuned the pre-trained DeBERTa encoder He et al.
+(2021) on these combined datasets.
+This model detects hallucinations based on the semantic relationship between the document and summary.
 Span-based detection
-NLI is a sentence level hallucination detection while NER is restricted to a predefined set of named entities, and we further explore more fine-grained hallucination detection.
-Specifically, we train a token level hallucination detection model to provide more fine-grained AI feedback, e.g.
-hallucinated text spans.
-We start with the dataset labeled for NLI model and further ask GPT4 to highlight the hallucinated text spans when it is hallucinated.
-As shown in Figure 3, we start with a pre-trained replaced token detection head from DeBERTa He et al.
-(2021), and adapt it with GPT-4 generated data like a sequence labeling task.
-We refer to this model as Span-based detection model (SBD).
-Figure 3: Span based hallucination detection
-Multi-source ensemble
-We believe the combination of multi-sources is necessary to further boost performance with minimal cost.
-For this part, we collected about 10k training examples from the target traffic, where the label (hallucinated vs not-hallucinated) is produced by GPT4 detection and the features are outputs from NER, NLI and SBD detection methods.
-We adopt the Gradient boosting decision tree to leverage the diverse AI feedback and finetune a model using scikit-learn to generate a single numerical value indicating whether the response or summary is hallucinated.
+NLI provides sentence- or summary-level hallucination detection, while NER is restricted to a predefined set of named entities.
+To explore a more general fine-grained hallucination detection, we train a token-level hallucination detection model to provide more detailed AI feedback, such as highlighting hallucinated text spans.
+Starting with the dataset labeled by GPT-4 for the NLI model, as described in Section 3.2, we further ask GPT-4 to highlight the hallucinated text spans if the text contains hallucinations.
+As shown in Figure 3, we initiate model training with a pre-trained Replaced Token Detection (RTD) head from DeBERTa He et al.
+(2021).
+We adapt this model using the GPT-4 generated data to determine if a token is part of a hallucinated span.
+We refer to this model as the Span-Based Detection model (SBD).
+Figure 3: Span-based hallucination detection
+4.2 Multi-Source Ensemble
+We believe that combining multiple sources is essential to further boost performance while minimizing costs.
+This approach is also convenient when we want to use a single score to control thresholds, allowing us to prioritize high precision during real implementation.
+For this part, we collected about 10,000 training examples, where the labels (hallucinated vs.
+not hallucinated) are produced by GPT-4 detection, and the features are entities and their confidence scores from the NER, confidence score from NLI, and the confidence score from SBD.
+We adopt the Gradient Boosting Decision Tree to leverage the diverse AI feedback and fine-tune a model using scikit-learn to generate a single numerical value indicating the confidence or likelihood that the text to be checked (i.e., response, summary, or single sentence) is hallucinated.
 4.3 GPT4-based rewriting
-With AI feedback from different sources, we compose a prompt based on a pre-defined template and ask GPT4 to fix it.
-Initially, we devise prompt version v1 that does an exhaustive chain-of-thoughts reasoning and does a full rewriting, which gives better performance but higher latency.
-Later on, to meet the latency requirement in the product, we developed rewriting prompt v2, which reduces chain-of-thought reasoning and then skip the rewriting if not hallucinated or otherwise only rewrite the hallucinated sentences instead of the full response.
+With AI feedback from different sources, we formulated prompts based on a pre-defined template to guide GPT-4 in correcting hallucinations. We explored two distinct rewriting prompts:
+Rewriting prompt v1.
+This does an exhaustive CoT reasoning or analysis to identify hallucinations in the text.
+It then does a complete rewriting to correct the hallucination while maintaining the coherence.
+Rewriting prompt v2.
+This prompt reduces the extent of CoT reasoning and opts to skip rewriting if no hallucinations are detected.
+When rewriting is necessary, it focuses solely on the hallucinated sentences rather than the entire text.
+This approach ensures only the essential changes are made to the original content.
 5 Main Results
 5.1 Detection Results
-We compare the different detection methods’ performance on both internal and public benchmark datset.
-As shown in Table 4, NER tends to have lower recall but relatively higher precision in public SummAC Laban et al.
+We compare the different detection methods’ performance on the following datasets:
+Internal benchmark dataset.
+This internal benchmark dataset has N=1400 examples consisting of 200 representative documents/transcripts x 7 systems of summaries.
+The ground-truth label is collected by our hired independent data vendors.
+Public benchmark dataset: SummAC.
+This dataset, as detailed in Laban et al.
+(2022) has N=1700 examples collected from six datasets focused on summary inconsistency detection, with ground-truth labels provided by humans in each datasets.
+We also reference the best results from that paper as baselines.
+We evaluated our detection methods against GPT-4 with a focus on both accuracy and latency.
+As shown in Table 4, NER tends to have lower recall but higher precision in public SummAC Laban et al.
 (2022).
-It’s notable that of the three tailored models, the SBD method outperforms the rest in all metrics, showcasing the effectiveness of detailed detection - an insight that is valuable to highlight.
-Table 4: Performance of different methods on the internal benchmark and public SummAC.
+It’s notable that of the three tailored models, the SBD method outperforms the rest in all metrics, showcasing the effectiveness of detection at the token level.
+Table 4: Performance of different methods on the internal benchmark and public SummAC. SBD methods is very competitive and worth further exploring.
+In Table 5, Compared with latency of GPT4 on the internal benchmark dataset, our finetuned NLI and SBD models enjoy significant latency advantages.
+Method Latency (s/request)
+NLI model 1.2s (on V100 GPU)
+SBD model 1.3s (on V100 GPU)
+GPT4 w/ CoT 7.9s (per API call)
+Table 5: Latency of finetuned models and the GPT4-based method.
 5.2 Mitigation Results
-We define mitigation rate as percentage of hallucination being fixed and use GPT4 output tokens length as proxy to the latency and cost.
-As shown in Table 5, prompt v2 is more token-efficient and achieves good balance of performance, latency and cost.
+Here is the evaluation setting:
+Benchmark dataset.
+This is an internal testset of 200 samples, derived from an application or feature team.
+This dataset includes 100 documents with two summaries from genearated GPT4 per document: one in paragraph format and one in bullet-point format.
+Metrics.
+We evaluated performance using two key metrics: mitigation rate measures the percentage of hallucination being successfully corrected fixed, as determined by GPT4-detection; and GPT-4 output tokens length is serving for as proxy metrics for evaluating the latency and cost.
+As shown in Table 6, Rewriting Prompt v2 is great token efficiency and achieves good a balanced trade-offce among rewriting quality, latency, and cost.
+Rewriting
+Prompt
+Mitigation
+Rate
+GPT-4 Output
+Tokens (avg)
+No rewrite 0.0 244 (original output)
+Prompt v1 66.0% 587
+Prompt v2 44.7% 130
 Table 6: Rewriting to balance the quality and latency.
 5.3 Performance in Production
-As mentioned above, we have developed two pipelines: 1) the detection-only pipeline and 2) the detection with mitigation pipeline.
-They are integrated into LLM-based products to mitigate the customer’s complaints about hallucination.
+We have developed two pipelines for production usage: 1) the detection-only pipeline and 2) the detection with mitigation pipeline.
+They have been integrated into LLM-based products to mitigate the customer’s complaints about hallucination as in Figure 4.
 In practice, we start with the detection-only pipeline and block the hallucinated context from affecting the customers.
-Gradually, we moved on to the full pipeline of detection and mitigation in the user experience as shown in Figure 4.
+Gradually, we move on to the full pipeline of detection and mitigation.
 Figure 4: User experience of Hallucination Detection & Mitigation Pipeline in Production
-Accurately measuring the effectiveness of hallucination detection and mitigation in a production environment is a complex task.
-We adopt two approaches: we conduct offline human evaluation on some benchmark datasets and we develop a mirror traffic-based flighting system that leverages GPT4 for online quality measurement.
-This online system applies mirrored traffic to various pipelines and uses GPT4 to assess hallucination rates and the overall quality of rewritten content, and it better monitors the dynamic nature of real-world applications.
-The human evaluation result in Table 6 shows that our detection pipelines are very effective.
-We experimented on key point generation datasets which contain 469 and 161 documents separately.
-On both datasets, the proposed methods are able to detect hallucinations with precision of above 50% and reduces more than half of the detected hallucinations with precision of above 60%.
-Table 6: Offline human evaluation of detection and mitigation.
-When applied to a real product scenario, the precision cannot be evaluated directly as we cannot check customer data.
-As shown by the online result in Table 7, the GPT4-based mitigation can greatly reduce the hallucinated responses and thus the block rate at expense of acceptable increase of latency.
-Table 7: Online monitoring and comparisons of the pipelines based on flighting traffic.
-6 Challenges, Learnings & Limitations
-We have encountered a number of considerable challenges when we integrate this federated hallucination detection and mitigation service hallucination into real products.
-We have suggested various workable solutions, although some are imperfect, and we outline them with the aim of assisting the community in acknowledging and addressing these challenges collaboratively.
+Accurately measuring the effectiveness of hallucination detection and mitigation in real-world applications is a complex task.
+To address this, we adopt two approaches: 1) offline human evaluation using a production-related benchmark dataset and 2) GPT4-based online monitoring based on the actual production traffic.
+For the offline human evaluation, we applied our detection and rewriting pipeline to a production-related testset consisting of bullet-point style summaries comprising of 630 individual keypoints.
+We have 25 (i.e.
+4.0% of 630) in total detected as containing hallucinations.
+An independent human data labeler to verified that 15 of the 25 dectections were accurate (i.e.
+check precision is 60.0%).
+Additionally, the human labeler to check if rewriting fixes the hallucinations, with 10 out 15 are accurately fixed (i.e.
+rewriting effective rate is 66.7%)
+In the online monitoring approach, we sampled a portion of production traffic - comprising pairs of <LLM input, LLM ouput > - for evaluation.
+GPT4 checked outputs for hallucinations and assessed if the rewriting process was necessary.
+Using GPT4’s judgement as a benchmark, we can observed that the precision of ensemble detection is above 80% (i.e.
+80% of the time is consistent with GPT4’s judgement) when the detection rate is about 3% and rewriting success rate is above 50%.
+In Table 7, we use statistics based on one-month production traffic to show both the detection-only pipeline and full pipeline of detection and mitigation can effectively reduce hallucinations, albeit an acceptable increase in latency.
+Table 7: Online monitoring and comparisons of the pipelines based on production traffic.
+6 Challenges and Future Work
 6.1 Measurement of Effectiveness in Production
-Accurately measuring the effectiveness of hallucination detection and mitigation and the values bringing to customers in a production environment is very challenging.
-We have designed a system which applies mirrored traffic to various pipelines and uses GPT4 to assess hallucination rates and the overall quality of rewritten content.
-However, the GPT4-based measurement has limitation and ensuring the reliability of these measurements and their alignment with human judgment remains an ongoing challenge, necessitating continuous refinement and validation.
+Accurately measuring the effectiveness of hallucination detection and mitigation, as well as the value they bring to customers in a production environment, is very challenging.
+We have designed a system that applies mirrored traffic to various pipelines and uses GPT4 to assess hallucination rates and the overall quality of rewritten content.
+However, the GPT4-based measurement has limitations and ensuring the reliability of these measurements and their alignment with human judgment remains an ongoing challenge, necessitating continuous refinement and validation.
 6.2 Handling Multilingual and Long Source Documents
-We have add several major non-English training data in NLI and SBD models to support the multilingual use case and using segmentation based approach to deal with long source documents.
-However, managing multilingual even cross-lingual inputs and long source documents adds another layer of complexity.
-Also it is still an open research problems how to do effective modeling of the long source document.
-6.3 Deep Customization Needs for Various Scenarios
-Different usages and user circumstances call for tailored adjustments.
-In the production, we’ve calibrated our ensemble-based detection for greater accuracy with a reduced block rate, while also adjusting our rewriting for decreased latency at the cost of some mitigation power.
-As we move forward, incorporating additional domain or task-specific datasets may enhance our detection capabilities.
-Concurrently, researching novel detection techniques or focusing exclusively on detecting certain types or severe cases of hallucinations might be more practical for real-world applications.
-Furthermore, we should consider investigating more efficient mitigation strategies involving smaller language models.
+We have incorporated major non-English training datasets into our NLI and SBD models to support multilingual use cases and are utilizing a segmentation-based approach to manage long source documents.
+However, handling inputs and outputs in different languages and their extensive combinations remains challenging.
+Additionally, developing effective models for processing long source documents continues to be an open research problem, requiring further exploration and innovation.
+6.3 Deep Customization Needs for Hallucination Handling
+Different user circumstances call for tailored adjustments to hallucination handling.
+For example, to meet the production needs, we’ve calibrated our ensemble-based detection for greater precision with a reduced block rate (or trigger rate) to avoid the availability issue, while also adjusting our rewriting for decreased latency at the cost of some mitigation power.
+However, there might be another setting, where we need a different balance of quality, latency and cost.
+Also, adapting to special domain or task or handling a specific types of hallucinations can also be great directions to explore.
 7 Conclusions
-To conclude, we developed a reliable and high-speed production service to detect and mitigate intrinsic or close-world hallucinations in LLMs.
-Our service combines named entity recognition (NER), natural language inference (NLI), span-based detection (SBD) models, and a decision tree analysis to identify diverse set of hallucinations with high accuracy.
-We also develop an efficient rewriting mechanism to ensure the quality and fidelity of LLM outputs.
-Hallucination is a key problem facing by LLM applications.
-Our journey from idea to deployment revealed several challenges in delivering high-quality solutions under real-world constraints.
-We had to balance latency, availability, and quality in production settings.
-Our work has broad applicability and potential for future impact, as LLMs are used in diverse sectors that require robust hallucination verification mechanisms.
-Our service can adapt to different LLM applications, each with different challenges and demands.
-A key next step is to address extrinsic or open-world hallucinations, which require integrating external knowledge bases and information verification systems.
+In this paper, we introduce a novel framework that can detect and mitigate intrinsic hallucinations, characterized by outputs not supported by grounding documents in LLMs.
+Our detection approach leverages the combined strengths of NER, NLI, and novel sequence labelling (SBD), and Decision Tree to detect as much as hallucination as possible.
+We further developed an effective LLM-based mitigation solution that blance the quality and latency.
+We detail the core elements of our framework and underscore the paramount challenges tied to response time, availability, and performance metrics, which are crucial for real-world deployment of these technologies.
+Our extensive evaluation, utilizing offline data and live production traffic, confirms the efficacy of our proposed framework and service.
 8 Ethical Considerations
@@
 Balancing these ethical imperatives with technical and operational demands is a complex but necessary challenge.
-Appendix A Appendix
-This is an appendix.